# Ferret: Federated Full-Parameter Tuning at Scale for Large Language Models

### v1

@@ 5 Experiments
In this section, we follow the practice in FedKSeed [12] to evaluate the efficacy of Ferret algorithm. We primarily compare Ferret with other federated full-parameter tuning baselines, including those based on zeroth-order methods, such as FedZO [20] and FedKSeed [12], as well as those based on first-order methods, such as FedAvg [15]. The evaluations are conducted using DataJuicer-1.3B [24] and LLaMA-3B [25] models on the Natural Instructions [26] and Dolly-15K [27] datasets. Since our Ferret is guaranteed to converge faster than zeroth-order FL methods, as shown in Sec. 4.2, we only run Ferret for 12 communication rounds compared to 40 rounds in FedKSeed on Natural Instructions, and 20 rounds versus 60 in FedKSeed on Dolly-15K. For additional details, please refer to Appx. C.1.
Table 2: Comparison of Rouge-L (%) among various algorithms. Each cell reports the mean $\pm$ std of Rouge-L scores from the final round of four runs, each using a different random seed.
@@
FedAvg 22.08 $\pm$ 1.52 27.88 $\pm$ 0.75 32.30 $\pm$ 1.23 34.27 $\pm$ 0.45
FedZO 21.74 $\pm$ 1.91 29.46 $\pm$ 0.38 32.91 $\pm$ 0.67 36.34 $\pm$ 0.39
FedKSeed 22.33 $\pm$ 1.72 29.77 $\pm$ 0.75 32.90 $\pm$ 0.37 35.64 $\pm$ 0.83
@@
Comparison on Accuracy. We present the model accuracy achieved by different federated tuning methods in Tab. 1. The results show that federated full-parameter tuning methods (including FedAvg, FedZO, FedKSeed, and Ferret) generally achieve better model accuracy compared to PEFT-based federated tuning methods (such as FedPTuning, FedPrompt, FedIT-SGD, and FedIT), underscoring the importance of full-parameter tuning for LLMs. Additionally, the results in Tab. 1 demonstrate that our proposed method, Ferret, consistently delivers strong or competitive performance across four different scenarios. E.g., on the Natural Instructions dataset, Ferret outperforms all other methods for different model sizes, with up to a 2.66% improvement over the next best method, FedKSeed. On the Dolly-15K dataset, Ferret maintains competitive performance. These results thus demonstrate the ability of our Ferret to sustain strong model accuracy in practice.
Table 3: Comparison of computational cost and communication overhead on LLaMA-3B, focusing on (a) the computational costs from local updates, global aggregation, and the overall tuning process; and (b) the per-round and overall communication costs. The improvement achieved by our Ferret is reported in brackets using blue (compared with FedKSeed) and orange (compared with FedAvg).
Algorithm Computational Cost (Sec.) Communication Cost (# param.)
Local Update Global Aggr. Overall Per-Round Overall
FedZO 58.1 4.8 2.5$\times$103 6.0$\times$109   2.4$\times$1011
FedKSeed 24.5 723.8 3.0$\times$104 8.2$\times$103 3.2$\times$105
FedAvg 2.3 4.8 85.2 6.0$\times$109   7.2$\times$1010
@@
Figure 1: Comparison of communication rounds required by Ferret and FedKSeed for convergence on Natural Instructions with (a) DataJuicer-1.3B and (b) LLaMA-3B.
@@ 6 Conclusion
In conclusion, our Ferret algorithm offers a highly desirable solution for the scalable, full-parameter tuning of LLMs in federated environments. By achieving high computational efficiency, fast convergence, and reduced communication overhead, Ferret overcomes the limitations of existing methods, striking an improved balance among these critical factors. Moreover, our rigorous theoretical analyses and extensive experiments validate Ferret as a robust and reliable approach for deploying LLMs in large-scale federated settings.
@@ Appendix A Related Work
Federated PEFT for LLMs. The field of federated learning (FL) has gained significant traction in its application to the fine-tuning of large language models (LLMs). Traditional FL approaches in this domain [8, 9, 28, 29] have predominantly focused on parameter-efficient fine-tuning (PEFT) techniques [5, 6, 30, 31, 32], which reduce the number of trainable parameters in LLMs to mitigate the extensive communication overheads in FL scenarios. Unfortunately, while PEFT methods such as those proposed in [9, 28] have shown promise, they often fall short in achieving the accuracy levels possible with full-parameter tuning [11], particularly in non-IID (non-independent and identically distributed) data settings commonly encountered in FL. In contrast, this paper focuses on federated full-parameter tuning of LLMs, aiming to achieve significantly reduced communication overhead while maintaining competitive model accuracy.
Federated Learning with Shared Randomness. Several approaches leveraging shared randomness have been proposed to enhance communication efficiency in FL. Methods including [12, 33, 34, 35, 36, 37, 38] demonstrate that by transmitting only a limited set of random seeds and scalar gradients, communication overhead can be drastically reduced. However, these methods rely on zeroth-order optimization (ZOO) for their local updates on each client. This reliance often results in poor scalability, as these methods require substantial computational costs per round to achieve the same local update progress and a larger number of communication rounds to converge compared with their first-order counterparts, such as FedAvg [15] and FedProx [16]. This limitation therefore becomes a bottleneck in large-scale federated environments. In contrast, our paper introduces the use of shared randomness within first-order FL, aiming to improve both computational and communication-round efficiency of zeroth-order FL. To the best of our knowledge, this is the first time that shared randomness has been introduced in first-order FL to reduce communication overhead.
@@ Appendix B Proofs
$\mathbb{E}\left[\widetilde{\Delta}\right]$ $=\frac{1}{\rho K}\mathbb{E}\left[{\mathbf{V}}{\mathbf{V}}^{\top}\right]\Delta$ (11)
$=\frac{1}{\rho K}\mathbb{E}\left[\sum_{k=1}^{K}{\mathbf{v}}_{k}{\mathbf{v}}_{k}^{\top}\right]\Delta$
$=\Delta\ .$
@@
Lemma 1 (Matrix Bernstein Inequality, Thm. 1.6.2 in [39]).
@@
$\left\|\sum_{k=1}^{K}\mathbb{E}\left[{\mathbf{X}}_{k}^{2}\right]\right\|$ $\leq\sum_{k=1}^{K}\left\|\frac{\rho}{K^{2}}{\mathbf{I}}_{d}\right\|$ (15)
$\leq\frac{\rho}{K}\ .$
@@
$\ell({\mathbf{w}}+\epsilon{\mathbf{v}}_{k};{\mathbf{x}}^{(i)})-\ell({\mathbf{w}};{\mathbf{x}}^{(i)})$ $\leq\epsilon\left(\nabla\ell({\mathbf{w}};{\mathbf{x}}^{(i)})\right)^{\top}{\mathbf{v}}_{k}+\frac{1}{2}\beta\epsilon^{2}\left\|{\mathbf{v}}_{k}\right\|^{2}$ (17)
$\leq\epsilon\left(\nabla\ell({\mathbf{w}};{\mathbf{x}}^{(i)})\right)^{\top}{\mathbf{v}}_{k}+\frac{1}{2}\beta\epsilon^{2}\ .$
@@
When $\epsilon\rightarrow 0$, (18) indicates that $g_{k}={\mathbf{v}}_{k}^{\top}\nabla\ell({\mathbf{w}};{\mathbf{x}}^{(i)})$, implying that this scalar gradient in zeroth-order method, e.g., FedKSeed [12], is an approximation of directional derivative, i.e., our projected update in (6) when $\Delta$ is replaced with $\nabla\ell({\mathbf{w}};{\mathbf{x}}^{(i)})$.
@@
$dK$ $=\left(\sum_{l\in[L]}d_{l}\right)\left(\sum_{l\in[L]}K_{l}\right)$
$=\sum_{l\in[L]}d_{l}\left(\sum_{l\in[L]}K_{l}\right)$
$>\sum_{l\in[L]}d_{l}K_{l}\ ,$
@@
Prop. 2 provides a looser bound than Thm. 2, primarily owing to the inequality $(b)$ in (20). Based on this looser bound, one might expect that block-wise reconstruction would incur a larger error compared to the vanilla reconstruction in (7). However, empirical results in Appx. C.4 and Appx. C.5 show that block-wise reconstruction yields comparable performance to the vanilla approach.
@@
Of note, we follow the general idea in [40] to prove the convergence of Ferret. To begin with, we introduce the following lemmas borrowed from [40]:
@@
where $(a)$ is from the local update of ${\mathbf{w}}_{r,t-1}^{\smash{(i)}}$ on each client $i$, and $(c)$ is based on (26) in Lemma 2 with $a=1/(T-1)$ and ${\mathcal{L}}({\mathbf{w}}_{r,t-1}^{(i)})={\mathcal{L}}^{(i)}({\mathbf{w}}_{r,t-1}^{(i)})$. Besides, $(d)$ results from Lemma 3 and the assumption that $\mathbb{E}[\|\nabla{\mathcal{L}}^{\smash{(i)}}({\mathbf{w}})-\ell({\mathbf{w}};{\mathbf{x}})\|^{2}]\leq\sigma^{2}$. Finally, $(e)$ comes from the summation of geometric series and the fact that $\eta\beta\leq 1/T$ as well as
@@
$\leq(19\eta^{2}T^{2}\beta-\eta T)\left\|\nabla{\mathcal{L}}({\mathbf{w}}_{r})\right\|^{2}+18\eta^{2}T^{2}\beta\sigma^{2}$
By summarizing both sides over $R$ rounds and letting $\eta\leq\frac{1}{20T\beta}$,
$\left\|\nabla{\mathcal{L}}({\mathbf{w}}_{r})\right\|^{2}\leq\frac{20\left({\mathcal{L}}({\mathbf{w}}_{r})-{\mathcal{L}}({\mathbf{w}}_{r+1})\right)}{\eta T}+360\eta T\beta\sigma^{2}\ ,$ (34)
Finally, we have the following results to conclude our proof:
$\min_{r\in[R)}\left\|\nabla{\mathcal{L}}({\mathbf{w}}_{r})\right\|^{2}\leq\frac{20\left({\mathcal{L}}({\mathbf{w}}_{0})-\min_{{\mathbf{w}}}{\mathcal{L}}({\mathbf{w}})\right)}{\eta TR}+360\eta\beta T\sigma^{2}\ .$ (35)

### v3

@@ 5 Experiments
In this section, we evaluate the efficacy of Ferret, following the practice in FedKSeed (Qin et al., 2024). We primarily compare Ferret with other federated full-parameter tuning baselines, including both zeroth-order methods (e.g., FedZO (Fang et al., 2022) and FedKSeed (Qin et al., 2024)) and first-order methods (e.g., FedAvg (McMahan et al., 2017)). Our evaluations use DataJuicer-1.3B (Chen et al., 2023) and LLaMA-3B (Touvron et al., 2023a) on the Natural Instructions (Wang et al., 2022) and Dolly-15K (Conover et al., 2023) datasets, as well as larger models (i.e., LLaMA2-7B and LLaMA2-13B (Touvron et al., 2023b)) on the CodeAlpaca (Chaudhary, 2023) and GSM8K (Cobbe et al., 2021).
FL Settings. In each round of federated learning, 5% of clients were randomly selected to participate. Following the same practice in FedKSeed (Qin et al., 2024), we set the total number of communication rounds to 40 for the NI dataset and 60 for Dolly-15K for all baselines. Due to the compelling efficiency of our method, we set the total number of communication rounds to 12 for the NI dataset and 20 for Dolly-15K for Ferret. However, for more complex tasks such as CodeAlpaca and GSM8K, we run all algorithms, including our Ferret, for 20 rounds to ensure a fair comparison. First-order baselines trained locally for one epoch, and FedKSeed trained for 200 steps, while our Ferret algorithm trained for 10 iterations (i.e., $T=10$ in Algo. 1). The $K$ value was set to 4096 for FedKSeed. All approaches perform local update with a batchsize of 1 to reduce memory consumption. For each local update iteration in Ferret, we accumulate the gradients from 4 samples.
More experimental details and ablation studies are provided in Appx. C.1 and Appx. C.5, C.6 respectively.
Table 2: Comparison of Rouge-L (%) among various algorithms. Each cell reports the mean $\pm$ std of Rouge-L scores from the final round of four runs, each with a different random seed. The highest and second-highest scores are shown in bold and underline, respectively.
@@
FedZO 21.74 $\pm$ 1.91 29.46 $\pm$ 0.38 26.99 $\pm$ 0.17 31.67 $\pm$ 0.35
FedKSeed 22.33 $\pm$ 1.72 29.77 $\pm$ 0.75 30.91 $\pm$ 0.29 34.56 $\pm$ 0.28
FedAvg 23.95 $\pm$ 2.76 32.11 $\pm$ 0.70 29.67 $\pm$ 1.26 30.98 $\pm$ 1.66
@@
Table 3: More comparison of Rouge-L (%) among various algorithms. Each cell reports the mean $\pm$ std of Rouge-L scores from the final round of four runs, each with a different random seed. The highest and second-highest scores are shown in bold and underline, respectively.
Algorithm CodeAlpaca GSM8K
LLaMA2-7B LLaMA2-13B LLaMA2-7B LLaMA2-13B
FedIT 4.66 $\pm$ 0.18 6.10 $\pm$ 0.18 30.31 $\pm$ 0.29 13.46 $\pm$ 0.34
FedZO 4.58 $\pm$ 0.26 6.19 $\pm$ 0.32 30.41 $\pm$ 0.31 13.63 $\pm$ 0.34
FedKSeed 8.33 $\pm$ 0.98 10.70 $\pm$ 0.47 28.26 $\pm$ 3.60 33.67 $\pm$ 1.15
FedAvg 15.41 $\pm$ 0.43 14.68 $\pm$ 0.26 38.30 $\pm$ 0.40 39.82 $\pm$ 0.17
Ferret (ours) 12.10 $\pm$ 0.47 11.84 $\pm$ 0.91 36.10 $\pm$ 1.18 34.50 $\pm$ 1.42
Table 4: Comparison of per-round computational cost and communication overhead on LLaMA-3B, including (a) the computational costs of local updates, global aggregation, and overall cost per round; and (b) the per-round communication cost. The improvement achieved by our Ferret is reported in brackets using blue (compared with FedKSeed) and orange (compared with FedAvg).
Algorithm Computational Cost (Sec.) Communication Cost
Local Update Global Aggr. Overall (# params.)
FedZO 32.6 0.3 32.9 6.0$\times$109
FedKSeed 56.9 123.8 180.7 8.2$\times$103
FedAvg 1.8 0.3 2.1 6.0$\times$109
[…]
Table 5: Comparison of per-round computational cost and communication overhead on LLaMA2-7B, focusing on (a) the computational costs from local updates, global aggregation, and the overall cost per round; and (b) the per-round communication cost. The improvement achieved by our Ferret is reported in brackets using blue (compared with FedKSeed) and orange (compared with FedAvg).
Algorithm Computational Cost (Sec.) Communication Cost
Local Update Global Aggr. Overall (# params.)
FedZO 54.1 0.7 54.8 1.4$\times$1010
FedKSeed 117.0 510.0 627.0 8.2$\times$103
FedAvg 5.8 0.7 6.5 1.4$\times$1010
@@
Figure 2: Comparison of communication rounds required by Ferret, FedKSeed, and FedAvg for convergence on Natural Instructions with (a) DataJuicer-1.3B and (b) LLaMA-3B.
5.1 Comparison on Accuracy
We present the model accuracy of different federated tuning methods in Tab. 2 and 3. The results in Tab. 2 demonstrate that federated full-parameter tuning methods (including FedAvg, FedZO, FedKSeed, and Ferret) generally achieve better model accuracy compared to PEFT-based federated tuning methods (such as FedPTuning, FedPrompt, FedIT-SGD, and FedIT). This underscores the importance of full-parameter tuning for Large Language Models (LLMs). Importantly, the results in both tables show that our proposed method consistently delivers strong or competitive performance across four different scenarios. Specifically, on the Natural Instructions dataset, our method outperforms all others for different model sizes, with up to a 2.66% improvement over the next best method, FedKSeed. On the Dolly-15K dataset, our method maintains competitive performance. Moreover, on both the CodeAlpaca and GSM8K datasets, our method achieves noticeably improved accuracy over FedIT and other zeroth-order baselines (i.e., FedZO and FedKseed). Of note, Ferret slightly underperform FedAvg, likely due to reconstruction errors caused by our method for these complex tasks. Overall, these results have well demonstrated the ability of our method to sustain strong model accuracy in practice across various datasets and model sizes.
5.2 Comparison on Scalability
@@
In Fig. 2, we also compare the convergence speeds of our Ferret with other baselines (e.g., FedKSeed and FedAvg) on Natural Instructions (with DataJuicer-1.3B and LLaMA-3B). The findings show that, Ferret converges remarkably fast, requiring only two communication rounds in line with FedAvg compared to the 40 rounds needed by FedKSeed. This results in a 20$\times$ reduction in communication round complexity for both DataJuicer-1.3B and LLaMA-3B.
5.3 Comparison on GPU Memory Consumption
Table 6: Peak GPU memory footprint of different methods.
Method DataJuicer-1.3B LLaMA-3B
FedAvg 9.9 GB 19.1 GB
FedKSeed 3.5 GB 7.8 GB
Ferret (ours) 9.9 GB 19.1 GB
We further provide the comparison on GPU memory consumption for different methods in Tab. 6. The results demonstrate that our proposed method Ferret maintains the same peak GPU memory footprint as FedAvg, requiring 9.9 GB and 19.1 GB for DataJuicer-1.3B and LLaMA-3B models, respectively, on the Natural Instructions dataset. Although FedKSeed, as a zeroth-order method, exhibits lower memory requirements, our Ferret delivers superior converged performance (refer to Sec. 5.1) and improved scalability (refer to Sec. 5.2) without introducing additional memory overhead compared to the baseline FedAvg method.
5.4 Ablation Studies
Refer to caption Refer to caption
(a) Convergence (b) Generalization
Figure 3: Convergence and generalization of our Ferret under varying $K$ on Natural Instructions with DataJuicer-1.3B where $2K_{0}$ corresponds to the communication cost of 7.8$\times$103 per round in Tab. 4.
Convergence and Generalization of Ferret under Varying $K$. In Fig. 3, we present the convergence and generalization of Ferret under varying $K$ on the Natural Instructions dataset with DataJuicer-1.3B, using the same experimental setup as described in Appx. C.1. Notably, Fig. 3 shows that: (a) a larger number of random bases (i.e., a larger $K_{0}$) generally leads to improved convergence, while the generalization performance remains comparable; (b) $2K_{0}$ already provides compelling convergence and generalization performance, and further increasing $K$ yields only marginal improvements in convergence; and (c) a slight decrease in generalization performance as $K$ increases is likely due to the reduced regularization effect from noisy gradients.
Refer to caption Refer to caption
(a) Convergence (b) Generalization
Figure 4: Convergence and generalization of Ferret under varying allocation scheme of $K$ for our block-wise reconstruction, in which $K_{l}\propto\left\|\Delta_{l}\right\|$ corresponds to our results in the previous sections.
Convergence and Generalization of Ferret under Varying Allocation of $K$. In Fig. 4, we present the convergence and generalization of Ferret under different allocation schemes for $K$ in our block-wise reconstruction, using the same experimental setup described in Appx. C.1, where $K_{l}\propto\left\|\Delta_{l}\right\|$ corresponds to our results in Sec. 5. Notably, Fig. 4 shows that Ferret achieves both faster convergence and improved generalization performance by following the best practices guided by Prop. 2. These findings therefore validate the significance and correctness of our Prop. 2.
Refer to caption Refer to caption
(a) Convergence (b) Generalization & Time Cost
Figure 5: Convergence, generalization, and projection time cost per round of Ferret under varying $L$ on Natural Instructions with DataJuicer-1.3B where $L=194$ is used in Tab. 2.
Convergence and Generalization of Ferret under Varying $L$. In Fig. 5, we present the convergence, generalization, and projection time cost of Ferret under varying block sizes $L$ on the Natural Instructions dataset with DataJuicer-1.3B, using the same experimental setup as described in Appx. C.1. Notably, Fig. 5 shows that increasing the number of blocks (i.e., a larger $L$) leads to improved convergence and reduced time cost for projection and reconstruction, while the generalization performance remains comparable. This improved convergence is likely due to the logarithmic term in the reconstruction error of our (7), as a larger number of blocks reduces the dimensionality of each block, thereby minimizing reconstruction error. In addition, the reduced time cost aligns with our analysis in Sec. 4.1 and the empirical results shown in Fig. 8(c), further highlighting the efficacy of our block-wise reconstruction method (8).
@@ 6 Conclusion
In this paper, we introduce Ferret, which offers a highly desirable solution for scalable federated full-parameter tuning of LLMs. By achieving high computational efficiency, fast convergence, and reduced communication overhead, Ferret overcomes the limitations of existing methods, striking an improved balance among these critical factors. Moreover, our rigorous theoretical analyses and extensive experiments validate Ferret as a robust and reliable approach for deploying LLMs in large-scale federated settings. While our work focuses on the performance and efficiency of Ferret, a critical area for future research is the in-depth privacy analysis of our method and exploring privacy-preserving mechanisms to ensure data security.
Acknowledgements
This research/project is supported by the National Research Foundation, Singapore under its National Large Language Models Funding Initiative (AISG Award No: AISG-NMLP-$2024$-$001$). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore. This research/project is supported by SAP and Singapore’s Economic Development Board under the Industrial Postgraduate Programme.
Impact Statement
This paper presents work whose goal is to advance the field of federated learning for large language models. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Related Work
Federated Parameter-Efficient Fine-Tuning (PEFT) for LLMs. Federated learning (FL) has become a crucial approach for training large language models (LLMs) while preserving data privacy. A significant body of work in this area has focused on parameter-efficient fine-tuning (PEFT) techniques (Hu et al., 2022; Wei et al., 2024; Lester et al., 2021; Lin et al., 2024; Hu et al., 2024; Lau et al., 2024) to address the substantial communication costs associated with training massive LLMs in distributed settings. These methods, such as those explored in (Zhang et al., 2023; Kuang et al., 2024; Zhang et al., 2024a; Kuang et al., 2023), aim to reduce the number of trainable parameters, thereby decreasing communication overhead. While PEFT approaches offer a viable solution, they often compromise on model accuracy compared to full-parameter fine-tuning, particularly in challenging non-IID (non-independent and identically distributed) data scenarios common in FL (Pu et al., 2023). In contrast, our work tackles the challenge of federated full-parameter tuning of LLMs, aiming to achieve accuracy levels comparable to centralized training without the prohibitive communication costs. This gap in achieving optimal accuracy with existing federated PEFT methods underscores the need for our approach.
Federated Learning with Gradient Compression. Another line of research explores gradient compression techniques to reduce communication overhead in federated learning. Methods like top-k sparsification (Lin et al., 2018), signSGD (Bernstein et al., 2018), and sketching-based compression (Rothchild et al., 2020) aim to reduce the size of transmitted gradients. Yet, these methods often achieve limited compression ratios and may not be effective for the high-dimensional parameter spaces of LLMs, particularly during full-parameter tuning. The millions of parameters involved in full-parameter LLM tuning far exceed the compression capabilities of these approaches. Furthermore, these techniques are not specifically tailored for the unique characteristics of LLMs, potentially limiting their effectiveness. In contrast, our proposed method leverages shared randomness within first-order FL to achieve significantly higher compression rates, scaling down the communication cost by orders of magnitude. This specialization for LLMs provides a critical advantage.
Federated Learning with Shared Randomness. The use of shared randomness has emerged as a promising avenue for reducing communication costs in FL. Methods like (Qin et al., 2024; Xu et al., 2023; Maritan et al., 2023; Feng et al., 2023; Dorfman et al., 2023; Zelikman et al., 2023; Rahimi et al., 2024) demonstrate that transmitting only random seeds and scalar gradients can drastically reduce communication overhead. However, these methods typically rely on zeroth-order optimization (ZOO) for local updates on each client. ZOO methods, while communication-efficient, are computationally expensive, requiring more rounds to converge than first-order methods like FedAvg (McMahan et al., 2017) and FedProx (Li et al., 2020). This computational inefficiency hinders scalability, particularly in large-scale federated environments. In contrast, our work introduces the use of shared randomness within first-order FL, combining the communication efficiency of shared randomness with the computational efficiency of first-order optimization. To the best of our knowledge, this is the first time that shared randomness has been successfully integrated with first-order FL for LLM fine-tuning, addressing a critical gap in the field.
@@ Appendix B Proofs
$\mathbb{E}\left[\widetilde{\Delta}\right]=\frac{1}{\rho K}\mathbb{E}\left[{\mathbf{V}}{\mathbf{V}}^{\top}\right]\Delta=\frac{1}{\rho K}\mathbb{E}\left[\sum_{k=1}^{K}{\mathbf{v}}_{k}{\mathbf{v}}_{k}^{\top}\right]\Delta=\Delta\ .$ (11)
@@
Lemma 1 (Matrix Bernstein Inequality, Thm. 1.6.2 in (Tropp et al., 2015)).
@@
$\left\|\sum_{k=1}^{K}\mathbb{E}\left[{\mathbf{X}}_{k}^{2}\right]\right\|\leq\sum_{k=1}^{K}\left\|\frac{\rho}{K^{2}}{\mathbf{I}}_{d}\right\|\leq\frac{\rho}{K}\ .$ (15)
@@
$\ell({\mathbf{w}}+\epsilon{\mathbf{v}}_{k};{\mathbf{x}}^{(i)})-\ell({\mathbf{w}};{\mathbf{x}}^{(i)})\leq\epsilon\left(\nabla\ell({\mathbf{w}};{\mathbf{x}}^{(i)})\right)^{\top}{\mathbf{v}}_{k}+\frac{1}{2}\beta\epsilon^{2}\left\|{\mathbf{v}}_{k}\right\|^{2}\leq\epsilon\left(\nabla\ell({\mathbf{w}};{\mathbf{x}}^{(i)})\right)^{\top}{\mathbf{v}}_{k}+\frac{1}{2}\beta\epsilon^{2}\ .$ (17)
@@
When $\epsilon\rightarrow 0$, (18) indicates that $g_{k}={\mathbf{v}}_{k}^{\top}\nabla\ell({\mathbf{w}};{\mathbf{x}}^{(i)})$, implying that this scalar gradient in zeroth-order method, e.g., FedKSeed (Qin et al., 2024), is an approximation of directional derivative, i.e., our projected update in (6) when $\Delta$ is replaced with $\nabla\ell({\mathbf{w}};{\mathbf{x}}^{(i)})$.
@@
$dK=\left(\sum_{l\in[L]}d_{l}\right)\left(\sum_{l\in[L]}K_{l}\right)=\sum_{l\in[L]}d_{l}\left(\sum_{l\in[L]}K_{l}\right)>\sum_{l\in[L]}d_{l}K_{l}\ ,$
@@
Prop. 2 provides a looser bound than Thm. 2, primarily owing to the inequality $(b)$ in (20). Based on this looser bound, one might expect that block-wise reconstruction would incur a larger error compared to the vanilla reconstruction in (7). However, empirical results in Appx. C.5 and Appx. C.6 show that block-wise reconstruction yields comparable performance to the vanilla approach.
@@
Of note, we follow the general idea in (Shu et al., 2024) to prove the convergence of Ferret. To begin with, we introduce the following lemmas borrowed from (Shu et al., 2024):
@@
where $(a)$ is from the local update of ${\mathbf{w}}_{r,t-1}^{\smash{(i)}}$ on each client $i$, and $(c)$ is based on (26) in Lemma 2 with $a=1/(T-1)$ and ${\mathcal{L}}({\mathbf{w}}_{r,t-1}^{(i)})={\mathcal{L}}^{(i)}({\mathbf{w}}_{r,t-1}^{(i)})$. Besides, $(d)$ results from Lemma 3 and the assumption that $\mathbb{E}[\|\nabla{\mathcal{L}}^{\smash{(i)}}({\mathbf{w}})-\nabla\ell({\mathbf{w}};{\mathbf{x}})\|^{2}]\leq\sigma^{2}$. Finally, $(e)$ comes from the summation of geometric series and the fact that $\eta\beta\leq 1/T$ as well as
@@
$\leq(19\eta^{2}T^{2}\beta-\eta T)\mathbb{E}\left[\left\|\nabla{\mathcal{L}}({\mathbf{w}}_{r})\right\|^{2}\right]+18\eta^{2}T^{2}\beta\sigma^{2}\ .$
By rearranging and letting $\eta\leq\frac{1}{20T\beta}$, we have
$\mathbb{E}\left[\left\|\nabla{\mathcal{L}}({\mathbf{w}}_{r})\right\|^{2}\right]\leq\frac{20\ \mathbb{E}\left[{\mathcal{L}}({\mathbf{w}}_{r})-{\mathcal{L}}({\mathbf{w}}_{r+1})\right]}{\eta T}+360\eta T\beta\sigma^{2}\ ,$ (34)
Finally, by summarizing both sides over $R$ rounds and scaling them with $1/R$, we have the following results to conclude our proof:
$\min_{r\in[R)}\mathbb{E}\left[\left\|\nabla{\mathcal{L}}({\mathbf{w}}_{r})\right\|^{2}\right]\leq\frac{20\left({\mathcal{L}}({\mathbf{w}}_{0})-\min_{{\mathbf{w}}}{\mathcal{L}}({\mathbf{w}})\right)}{\eta TR}+360\eta\beta T\sigma^{2}\ .$ (35)